OpenStudy (anonymous):

There are 25 more women than men in the night school Louise attends.Altogether there are 115 students in the school.How many women attend the night school?/

OpenStudy (anonymous):

try doing 115-25

OpenStudy (aripotta):

x = # of men x + (x + 25) = 115 2x + 25 = 115 2x = 90 x = 45 45 + 25 = 70 so there's 70 women